OpenStudy (anonymous):

need help: the area of a rectangle is 20x*squared*−27x−8 The length is 4x+1 What is the width?

OpenStudy (anonymous):

20x^2 - 27x - 8 = (4x + 1)(5x - 8) If this is an 8th grade problem, chances are that the length is 4x + 1, not 4x - 1. That way, the problem would come out neatly: the width would be 5x - 8. Can you check this?

OpenStudy (anonymous):

let's factor 20x^2-27x-8= We need factors of -160 which add to -27 we know we have 4x and 1 -32 and +5 factor 20 5 and 4 and -8 factor -32 4 and -8 -32 +5=-27 -32 4 and -8 (4x+1)*(5x-8) The width is 5x-8 We could also have divided 20x^2-27x-8 4x-1

OpenStudy (anonymous):

ok, thanks. So the answer is 4x - 1? PS this is an eight grade question

OpenStudy (anonymous):

no it is 5x - 1

OpenStudy (anonymous):

The width is 5x-8

OpenStudy (anonymous):

oh, ok. thanks

OpenStudy (anonymous):

welcome